# A Systematic Survey of Prompt Engineering in Large Language Models: Techniques and Applications

### v1

@@ 2.3.1 Retrieval Augmented Generation (RAG)
LLMs have revolutionized text generation, yet their reliance on limited, static training data hinders accurate responses, especially in tasks demanding external knowledge. Traditional prompting falls short, requiring expensive retraining. Retrieval Augmented Generation (RAG) Lewis et al. (2020) emerges as a novel solution, seamlessly weaving information retrieval into the prompting process. RAG analyzes user input, crafts a targeted query, and scours a pre-built knowledge base for relevant resources. Retrieved snippets are incorporated into the original prompt, enriching it with contextual background. The augmented prompt empowers the LLM to generate creative, factually accurate responses. RAG’s agility overcomes static limitations, making it a game-changer for tasks requiring up-to-date knowledge. RAG outperformed seq2seq models and task-specific architectures on ODQA benchmarks, achieving exact match scores, reaching up to 56.8% on TriviaQA and 44.5% on Natural Questions.
@@ 2.3.2 ReAct Prompting
Unlike previous studies that treated reasoning and action separately, ReAct Yao et al. (2022) enables LLMs to generate reasoning traces and task-specific actions concurrently. This interleaved process enhances synergy between reasoning and action, facilitating the model in inducing, tracking, and updating action plans while handling exceptions. ReAct is applied to diverse language and decision-making tasks, showcasing its effectiveness over state-of-the-art baselines. Notably, in question answering (HotpotQA) and fact verification (Fever), ReAct addresses hallucination and error propagation issues by interacting with a simple Wikipedia API, producing more interpretable task-solving trajectories. Additionally, in interactive decision-making benchmarks like ALFWorld and WebShop, ReAct surpasses both imitation and reinforcement learning approaches, achieving notable success rates of 34% and 10%, respectively, with minimal in-context examples.
@@ 2.3.3 Chain-of-Verification (CoVe) Prompting
To address hallucinations in LLMs, Dhuliawala et al. (2023) proposed Chain-of-Verification (CoVe), which involves a systematic four-step process including the model generate baseline responses, plan verification questions to check its work, answer the questions independently, and produce a revised response incorporating the verification. By verifying its work through this deliberate multi-step approach, the LLM enhances logical reasoning abilities and reduces errors even with contradictory information. CoVe emulates human verification to bolster the coherence and precision of LLM output. Experiments on list questions, QA, and long-form generation demonstrate that CoVe decreases hallucinations while maintaining facts. Focused verification questions help models identify and correct their inaccuracies.
@@ 2.3.5 Chain-of-Knowledge (CoK) Prompting
Traditional prompting techniques for LLMs have proven powerful in tackling basic tasks. However, their efficacy diminishes due to complex reasoning challenges, often resulting in unreliable outputs plagued by factual hallucinations and opaque thought processes. This limitation arises from their reliance on fixed knowledge sources, ineffective structured query generation, and lack of progressive correction that fails to guide the LLM adequately. Motivated by human problem-solving, CoK Li et al. (2023d) systematically breaks down intricate tasks into well-coordinated steps. The process initiates with a comprehensive reasoning preparation stage, where the context is established, and the problem is framed. Subsequently, it engages in a dynamic knowledge adaptation phase, meticulously gathering evidence from various sources, such as its internal knowledge base, external databases, and the given prompt.
@@ 2.7.1 Contrastive Chain-of-Thought (CCoT) Prompting
Traditional CoT prompting for LLMs often misses a crucial element: learning from mistakes. That is where Contrastive Chain-of-Thought Prompting (CCoT) Chia et al. (2023) dives in, providing both valid and invalid reasoning demonstrations alongside original prompts. Imagine exploring a map with the right path and the wrong turns to avoid – that is the advantage of contrastive CoT! This dual-perspective approach, tested on reasoning benchmarks like SQuAD and COPA, pushes LLMs to step-by-step reasoning, leading to 4-16% improvements in strategic and mathematical reasoning evaluations compared to traditional CoT, further improved by approximately 5% when integrated with self-consistency techniques. However, questions remain about this technique, such as the automated generation of contrasting demonstrations for diverse problems and its applicability to other NLP tasks beyond reasoning.
@@ 3 Conclusion
In the domain of artificial intelligence, prompt engineering has become a transformative force, unlocking the vast potential of LLMs. This survey paper aims to serve as a foundational resource that systematically categorizes 29 distinct prompt engineering techniques based on their targeted functionalities, inspiring further research and empowering innovators in the evolving landscape of prompt engineering. The analysis spans applications, models, and datasets, shedding light on the strengths and limitations of each approach. Furthermore, we have added a diagram and a table to highlight the important points. Despite the remarkable successes, challenges persist, including biases, factual inaccuracies, and interpretability gaps, necessitating further investigation and mitigation strategies. The future of prompt engineering holds immense potential, with emerging trends like meta-learning and hybrid prompting architectures promising amplified capabilities. However, ethical considerations are paramount, emphasizing responsible development and deployment to ensure positive integration into our lives.

### v2

@@ 2.3.1 Retrieval Augmented Generation (RAG)
LLMs have revolutionized text generation, yet their reliance on limited, static training data hinders accurate responses, especially in tasks demanding external knowledge. Traditional prompting falls short, requiring expensive retraining. Retrieval Augmented Generation (RAG) (Lewis et al., 2020) emerges as a novel solution, seamlessly weaving information retrieval into the prompting process. RAG analyzes user input, crafts a targeted query, and scours a pre-built knowledge base for relevant resources. Retrieved snippets are incorporated into the original prompt, enriching it with contextual background. The augmented prompt empowers the LLM to generate creative, factually accurate responses. RAG’s agility overcomes static limitations, making it a game-changer for tasks requiring up-to-date knowledge. RAG outperformed seq2seq models and task-specific architectures on ODQA benchmarks, achieving exact match scores, reaching up to 56.8% on TriviaQA and 44.5% on Natural Questions.
@@ 2.3.2 ReAct Prompting
Unlike previous studies that treated reasoning and action separately, ReAct (Yao et al., 2022) enables LLMs to generate reasoning traces and task-specific actions concurrently. This interleaved process enhances synergy between reasoning and action, facilitating the model in inducing, tracking, and updating action plans while handling exceptions. ReAct is applied to diverse language and decision-making tasks, showcasing its effectiveness over state-of-the-art baselines. Notably, in question answering (HotpotQA) and fact verification (Fever), ReAct addresses hallucination and error propagation issues by interacting with a simple Wikipedia API, producing more interpretable task-solving trajectories. Additionally, in interactive decision-making benchmarks like ALFWorld and WebShop, ReAct surpasses both imitation and reinforcement learning approaches, achieving notable success rates of 34% and 10%, respectively, with minimal in-context examples.
@@ 2.3.3 Chain-of-Verification (CoVe) Prompting
To address hallucinations in LLMs, Dhuliawala et al. (2023) proposed Chain-of-Verification (CoVe), which involves a systematic four-step process including the model generate baseline responses, plan verification questions to check its work, answer the questions independently, and produce a revised response incorporating the verification. By verifying its work through this deliberate multi-step approach, the LLM enhances logical reasoning abilities and reduces errors even with contradictory information. CoVe emulates human verification to bolster the coherence and precision of LLM output. Experiments on list questions, QA, and long-form generation demonstrate that CoVe decreases hallucinations while maintaining facts (Sahoo et al., 2024). Focused verification questions help models identify and correct their inaccuracies.
@@ 2.3.5 Chain-of-Knowledge (CoK) Prompting
Traditional prompting techniques for LLMs have proven powerful in tackling basic tasks. However, their efficacy diminishes due to complex reasoning challenges, often resulting in unreliable outputs plagued by factual hallucinations and opaque thought processes. This limitation arises from their reliance on fixed knowledge sources, ineffective structured query generation, and lack of progressive correction that fails to guide the LLM adequately. Motivated by human problem-solving, CoK (Li et al., 2023d) systematically breaks down intricate tasks into well-coordinated steps. The process initiates with a comprehensive reasoning preparation stage, where the context is established, and the problem is framed. Subsequently, it engages in a dynamic knowledge adaptation phase, meticulously gathering evidence from various sources, such as its internal knowledge base, external databases, and the given prompt.
@@ 2.7.1 Contrastive Chain-of-Thought (CCoT) Prompting
Traditional CoT prompting for LLMs often misses a crucial element: learning from mistakes. That is where Contrastive Chain-of-Thought Prompting (CCoT) (Chia et al., 2023) dives in, providing both valid and invalid reasoning demonstrations alongside original prompts. Imagine exploring a map with the right path and the wrong turns to avoid – that is the advantage of contrastive CoT! This dual-perspective approach, tested on reasoning benchmarks like SQuAD and COPA, pushes LLMs to step-by-step reasoning, leading to 4-16% improvements in strategic and mathematical reasoning evaluations compared to traditional CoT, further improved by approximately 5% when integrated with self-consistency techniques. However, questions remain about this technique, such as the automated generation of contrasting demonstrations for diverse problems and its applicability to other NLP tasks beyond reasoning.
@@ 2.9.4 Chain-of-Code (CoC) Prompting
Reasoning and Logic
Self-Refine Manual Multi GPT-3.5,GPT-4 7 diverse tasks(e.g., Dialogue Response,Math Reasoning) Task-specific (Accuracy, Human Preference)
Code Prompting LM Generated Multi GPT 3.5, Mixtral CondQA,ShaRC,BGQA F1
ECHO Hybrid Multi gpt-3.5-Turbo-0301 Arithmetic,Commonsense,Symbolic Accuracy
Logic-of-thought LM Generated Multi GPT 3.5-turbo, GPT-4 ReClor, LogiQA, RuleTaker, ProofWriter, FOLIO Accuracy
IAP Manual Multi LLaMA-3-8B-Instruct, Qwen-14B-Chat Math,Logic,Commonsense Accuracy
EEDP Manual Single GPT-4-turbo Merged 1000, ZINC test 2500 Accuracy
LoT LM Generated Multi GPT-4o Japanese Civil Law,Normative sentence Precision,Recall,F2
NoT LM Generated Single GPT-3.5,GPT-4, Mistral-7B,LLaMA3-8B ProScript,Schema-11,WikiHow Script F1,GED
BoT LM Generated Multi Llama3-8B, Llama3-70B 10 reasoning-intensive tasks (e.g., Game of 24, Geometric Shapes) Accuracy
CD-CoT Manual Single gpt-3.5-turbo-0613, Gemini-Pro(and others) Multiple tasks (e.g. BIG-Bench subsets, commonsense QA, etc.) Accuracy, Solve Rate, Human Preference
R-CoT Manual Single GPT4o, R-CoT-8B(and others) GeoMM,MathVista,GeoQA Accuracy
CoD Hybrid Single GPT-4o,Claude 3.5 Sonnet Arithmetic, Commonsense, Symbolic Accuracy
@@ 3 Conclusion
In the domain of artificial intelligence, prompt engineering has become a transformative force, unlocking the vast potential of LLMs. This survey paper aims to serve as a foundational resource that systematically categorizes 41 distinct prompt engineering techniques based on their targeted functionalities, inspiring further research and empowering innovators in the evolving landscape of prompt engineering. The analysis spans applications, models, and datasets, shedding light on the strengths and limitations of each approach. Furthermore, we have added a diagram and a table to highlight the important points. Despite the remarkable successes, challenges persist, including biases, factual inaccuracies, and interpretability gaps, necessitating further investigation and mitigation strategies. The future of prompt engineering holds immense potential, with emerging trends like meta-learning and hybrid prompting architectures promising amplified capabilities. However, ethical considerations are paramount, emphasizing responsible development and deployment to ensure positive integration into our lives.